Shelia invested 25,000 in two funds A and B. Fund A has a yield of 5%, while fund B has a yield of 4%. If the income for one year is $1,140, how much money was invested in each fund?
Answer by mananth(16949)

let her invest $x in fund A 5%
So she will invest the balance 25000-x in fund B 4%
Yield = $1140
0.05x +0.04(25000-x) = 1140
0.05x +1000.00- 0.04x= 1140
0.01x = 1140 - 1000
0.01x = 40
x+40/0.01
= $ 4000 in fund A
$21000 in fund B